# AboutMe: Using Self-Descriptions in Webpages to Document the Effects of English Pretraining Data Filters

### v1

@@ Abstract
Large language models’ (LLMs) abilities are drawn from their pretraining data, and model development begins with data curation. However, decisions around what data is retained or removed during this initial stage is under-scrutinized. In our work, we ground web text, which is a popular pretraining data source, to its social and geographic contexts. We create a new dataset of 10.3 million self-descriptions of website creators, and extract information about who they are and where they are from: their topical interests, social roles, and geographic affiliations. Then, we conduct the first study investigating how ten “quality” and English language identification (langID) filters affect webpages that vary along these social dimensions. Our experiments illuminate a range of implicit preferences in data curation: we show that some quality classifiers act like topical domain filters, and langID can overlook English content from some regions of the world. Overall, we hope that our work will encourage a new line of research on pretraining data curation practices and its social implications.
[Uncaptioned image] Code
github.com/lucy3/whos_filtered
@@
Li Lucy1,2   Suchin Gururangan5   Luca Soldaini1 Emma Strubell1,4David Bamman2Lauren Klein3 Jesse Dodge1 1Allen Institute for AI   2University of California, Berkeley  3Emory University 4Carnegie Mellon University  5University of Washington lucy3_li@berkeley.edu
@@ 1 Introduction
Large language models (LLMs) are increasingly being incorporated into real-world applications. These LLMs are sometimes described to be “general-purpose,” and some suggested uses include simulating human behaviors Markel et al. (2023); Argyle et al. (2023); Aher et al. (2023). However, studies around these models’ downstream behavior have shown that they reflect a limited and skewed set of human knowledge and perspectives Johnson et al. (2022); Durmus et al. (2023); Atari et al. (2023).
@@
Past work has shown that the composition of pretraining data impacts model behavior Kandpal et al. (2023); Razeghi et al. (2022); Chang et al. (2023); Gonen et al. (2023). Therefore, documentation of this data facilitates informed and appropriate application of models Gebru et al. (2021). We argue that it is additionally important to examine how data is transformed prior to pretraining, and document the implications of transformation steps. Pretraining data curation involves many decision points, which may be motivated by performance on popular benchmarks (e.g. Rae et al., 2021), or simply by some notion of text “quality.” There remains many under-examined assumptions within data curation pipelines, and our paper investigates how common filtering practices may impact web content from various sources.
Figure 1: A paraphrased excerpt from a website’s about page, with commonly stated social dimensions emphasized with highlighting.
@@
We provide a new dataset and framework, AboutMe, for documenting data filtering’s effects on text grounded to social and geographic contexts. Sociolinguistic analyses in NLP are limited by a lack of large-scale, self-reported sociodemographic information tied to language data Holstein et al. (2019); Andrus et al. (2021). Though text can be attributed to broad sources, such as Wikipedia, the backgrounds of content creators at more granular levels are often unknown. In particular, web crawl data lacks consistent and substantive user metadata. In our study, we leverage existing structure found in web data. Specifically, some websites include pages delineated to be about the website creator, such as an “about me” page (Figure 1). Thus, we are able to identify whose language is represented in web scraped text at an unprecendented scale.
@@
From websites’ about pages, we measure their topical interests, their positioning as individuals or organizations, their self-identified social roles, and their associated geographic locations (§ 2). We then apply ten “quality” and English ID filters drawn from prior literature on LLM development (§ 3) onto these websites, show whose pages are removed or retained, and investigate possible reasons for filters’ preferences (§ 4). Together, our experiments uncover behavioral patterns within and across filters tied to aspects of websites’ provenance. We find that model-based “quality” filters’ implicit preferences for certain topical domains lead to text specific to different roles and occupations being removed at varying rates. In addition, English content associated with non-anglophone regions of the world can be removed due to filtering approaches that assume pages are monolingual. To facilitate future work, we release our code, data, and other resources, which will be available on Github.
@@ 2 Extracting Social Dimensions from about Pages
The field of sociolinguistics conceptualizes language as a performance of one’s social identity, or membership in a social group Nguyen et al. (2016). Websites’ about pages capture aspects of their creators’ social identities that they deem salient and significant enough to mention in a summary (Figure 1). Thus, these self-descriptions can help delineate meaningful differences in language varieties and use. We extract social aspects that are present across large sets of pages using automated methods, some of which we contribute as novel approaches. We do not examine attributes such as race or gender, as these are less commonly explicitly stated on about pages and may raise the risk of mismeasurement (§ 7).
2.1 Data preprocessing
AboutMe is derived from twenty four public snapshots of Common Crawl collected between 2020–05 to 2023–06. We extract text using CCNet Wenzek et al. (2020) and deduplicate URLs across all snapshots. Our study focuses on data curation of LLMs that prioritize English or intend to be monolingually English, and our pipeline for identifying social aspects of websites uses methods that work best for English. Thus, we limit our study to the English subset of the CCNet output, which are webpages that have a fastText English ID score $>0.5$.
From this Common Crawl data, we identify website hostnames that include an about page, which we define as URL paths containing about, about-me, about-us, or bio (Appendix A). We then pair each about page with a random page on the same website. AboutMe thus contains both information about the creator/s of a website and a sample of their textual content (Table 1). Though we use this dataset to study LLM data curation practices, text linked to descriptions of their creators can also facilitate other areas of research, including self-presentation Sun et al. (2023); Pathak et al. (2021) and language variation Nguyen et al. (2016).
2.2 Topical interests
@@
First, we treat about pages as summaries of website creators’ interests and topical focus. Following past work on the unsupervised discovery of text domains Gururangan et al. (2023), we embed about pages using unigram counts and tf-idf Manning et al. (2008) and cluster them with balanced $k$-means. We set $k=50$ and surface a wide range of topical clusters in our data, including design, finance, food, religion, and travel (Figure 2, Appendix C). Thus, this clustering step provides an initial broad overview of who and what is in AboutMe.
@@
Websites are either created by individuals, such as bloggers, or by organizations, such as businesses. We deem webpages with urls that contain about-me or bio as individuals, while those that are labeled as about-us are organizations. However, some pages are ambiguously labeled with about, and so we classify these into individuals or organizations by training a binary random forest classifier on labeled pages. Classifier inputs include several count features that are agnostic to pages’ topical content: the proportion of words on an about page in common pronoun series (he, she, they, we, I), the proportion of words that are tagged as a person by spaCy named-entity recognition, and the number of unique person first tokens. Our classifier achieves an average macro F1 of 89.2, via 5-fold cross validation on 10k examples per class. Hyperparameter choices, classifier confidence, and other implementation details are in Appendix D.1. By applying this classifier, we find that three-fourths of hostnames are organizations rather than individuals (Table 1).
@@ 2.4 Social roles
String-matching can be imprecise due to polysemy and mentions of people on about pages could refer to others (e.g. a customer). Thus, our role extraction approach targets explicit expressions of self-identification (e.g. I am a designer, entrepreneur, and mother). We hand-label a sample of 1K about page sentences that contain a diverse set of potential roles, [^1] https://en.wiktionary.org/w/index.php?title=Category:en:People and treat role extraction as a binary, sentence-level token classification task. Our full criteria for role annotation can be found in Appendix F.1, and our annotators achieved high agreement (Cohen’s $\kappa$ = 0.836).
@@
With this approach, we are able to identify social roles on 77.7% of all individuals’ about pages (Table 1). Pages that have any roles contain 5.5 on average (SD = 9.9), and Appendix F.2 lists the top 100 most frequently retrieved terms. When possible, we group terms into occupations based on a taxonomy created by the Occupational Information Network, or O*NET, e.g. an attorney and a lawyer are in the occupation of Lawyer in the Legal occupation family (Table 2, Appendix F.3). For our filtering rate analysis (§ 4), we include 780 social roles and occupations that occur at least 1K times in AboutMe.
@@
Arts, Design, Entertainment, Sports, & Media
1.1M
artist, director, designer, writer, photographer, musician, player
Production
620K
designer, engineer, maker, builder, operator, mechanic
Community & Social Service
452K
therapist, educator, advisor, pastor, activist, social worker
Computer & Mathematical
365K
engineer, developer, scientist, strategist, programmer
Educational Instruction & Library
308K
teacher, professor, lecturer, curator, tutor, graduate student
Table 2: Five most common occupation families in AboutMe, by website count, with example social roles. An extended version of this table is in Appendix F.3.
@@ 2.5 Geography
We geoparse locations on about pages using Mordecai3, which tags named locations, retrieves candidate matches from a Geonames index, and disambiguates them using textual context Halterman (2023). For example, I’m from Alexandria, Virginia would be geoparsed to a location in the United States instead of Egypt. Mordecai3 is free and uses a local index, and so it can be scaled up to millions of webpages. With this approach, we are able to link 63.2% of websites in AboutMe to a country. This coverage exceeds the 10.38% obtained by using country-specific top-level domains, e.g. .uk for the United Kingdom Cook and Brinton (2017).
Locations mentioned on a page are often associated in some way with the creator/s of a website, but the strength of this association can vary. For example, a company may have been founded in one country, but ships products to another. Through manual annotation of 200 about pages, we find that 79.46% of websites with geoparsed countries originate from or reside in the most frequently referenced country on their about pages (evaluation details in Appendix G.1). Thus, we label each website with its most frequent country, yielding an overall website-level labeling accuracy of 91.0%. Due to the current global digital divide and our focus on English, the majority of websites in AboutMe are labeled with the United States and United Kingdom, with a long tail of other countries (Figure 3). For analysis, we group countries into 5 continental regions and 15 subregions delineated by the United Nations (Appendix G.2).
@@ 3 Pretraining Data Filters
Raw data is transformed into pretraining data for LLMs through a variety of steps (Penedo et al., 2023; Soldaini et al., 2023, inter alia), including deduplication Lee et al. (2022), decontamination (e.g. Touvron et al., 2023b), and explicit content filtering (e.g. OpenAI, 2023). Our study focuses on “quality” filtering and English langID. We select these two types of filters due to the subjectivity of how quality should be defined, and uncertainty around whether langID is robust to a wide range of language varieties Seargeant and Tagg (2011); Caswell et al. (2020).
Since web text may contain noisy content Eisenstein (2013), removing or downsampling “low quality” text is common in LLM development (Table 3). However, mismatch between filtering outcomes and downstream objectives can lead to performance degradation on some tasks Gao (2021); Longpre et al. (2023). In addition, Gururangan et al. (2022) use a dataset of high school newspapers to show that text from wealthier, more educated, and urban areas are less likely to be considered high quality by GPT-3’s model-based quality filter. We also critically examine the effects of quality filtering, but on a larger scale and across a variety of filters.
Figure 3: Common continental subregions in AboutMe. The most frequent countries are the United States, United Kingdom, India, Canada, Australia, China, Germany, New Zealand, Italy, and South Africa (see Appendix G for an extended list).
LangID is also a common step in data curation pipelines. Many LLMs are designed to focus only on English, and language-specific models use langID at the document-level in initial preprocessing stages (Table 3). LangID is also used to measure and adjust the composition of pretraining data for multilingual models Xue et al. (2021). However, popular langID systems are imperfect, due to reasons such as training and application domain mismatch and confusion between similar languages Kreutzer et al. (2022); Caswell et al. (2020).
We examine document-level quality and English filters, outlined in the following paragraphs, that are sufficiently documented in prior work (Table 3). Further details on the reproduction of each filter can be found in Appendix B. Descriptions of pretraining data curation are sometimes too vague or non-existent to allow for exact replication OpenAI (2023), but multiple recent and prominent LLMs still allude to the use of model- and heuristic-based data filters Touvron et al. (2023a); Gemini Team et al. (2023); Chowdhery et al. (2023).
@@
[…]
@@
Filter Examples of prior use Removal strategy
[…]
GPT-3 Brown et al. (2020)
Sampling based on scores
[…]
the Pile Gao et al. (2020)
Sampling based on scores
[…]
LLaMA Touvron et al. (2023a) & RedPajama Computer (2023)
Sampling based on scores
[…]
Specified in reference mixes by Xie et al. (2023), PaLM Chowdhery et al. (2023), and GPT-3 Brown et al. (2020)
Sampling based on scores
[…]
CCNet Wenzek et al. (2020)
Percentile cutoffs: 33.3% or 66.7%
[…]
Gopher Rae et al. (2021), Chinchilla Hoffmann et al. (2022), & RefinedWeb Penedo et al. (2023)
Specific cutoffs for each rule
[…]
CCNet Wenzek et al. (2020), LLaMA Touvron et al. (2023a), RefinedWeb Penedo et al. (2023)
Cutoffs: 0.50 (CCNet, LLaMA), 0.65 (RefinedWeb)
[…]
The Pile Gao et al. (2020)
Cutoff: 0.50
[…]
multilingual C4 Xue et al. (2021)
Cutoff: 0.70
[…]
C4 Dodge et al. (2021); Raffel et al. (2023)
Cutoff: 0.99
[…]
@@
[…]
@@ 4 Whose websites are filtered?
In this section, we overview the effects of data filters on sampled webpages grouped by social dimensions identified from their about pages. As discussed in § 3, definitions of “quality” differ across LLMs, and algorithmic measurements of “English” can be flawed. Broadly, we examine the degree of consensus among filters when scoring pages, and identify themes that characterize their behavior. Within some dimensions, we also investigate whether filtering rates reflect systemic differences in power and status among social groups Blank (2013); Davis (2018).
Past LLMs have chosen a range of score cutoffs and sampling mechanisms to reduce undesirable text (Table 3). With this variation in mind, we examine the outcome of model-based filters through the lens of two contrasting scenarios. First, whose pages are least affected, or retained, if we were to keep only the documents within a top percentile of scores? Second, whose pages are most affected, or removed, if we were to filter those at a very bottom percentile? We select top and bottom cutoff percentiles of 10% and 90%, though for CLD2 and langdetect, a large number of score ties meant that that cutoffs for both scenarios were 5.2% and 8.7%, respectively. For rule-based filters, we use cutoffs specified by Gopher. All cutoffs are listed in Appendix B.
@@
[…]

### v3

@@ Abstract
Large language models’ (LLMs) abilities are drawn from their pretraining data, and model development begins with data curation. However, decisions around what data is retained or removed during this initial stage are under-scrutinized. In our work, we ground web text, which is a popular pretraining data source, to its social and geographic contexts. We create a new dataset of 10.3 million self-descriptions of website creators, and extract information about who they are and where they are from: their topical interests, social roles, and geographic affiliations. Then, we conduct the first study investigating how ten “quality” and English language identification (langID) filters affect webpages that vary along these social dimensions. Our experiments illuminate a range of implicit preferences in data curation: we show that some quality classifiers act like topical domain filters, and langID can overlook English content from some regions of the world. Overall, we hope that our work will encourage a new line of research on pretraining data curation practices and its social implications.
@@
Li Lucy1,2   Suchin Gururangan5   Luca Soldaini1 Emma Strubell1,4David Bamman2Lauren F. Klein3 Jesse Dodge1 1Allen Institute for AI   2University of California, Berkeley  3Emory University 4Carnegie Mellon University  5University of Washington lucy3_li@berkeley.edu
@@ 1 Introduction
Large language models (LLMs) are sometimes described to be general-purpose (e.g. Radford et al., 2019), and are increasingly incorporated into real-world applications. However, their behavior can reflect a limited set of human knowledge and perspectives Johnson et al. (2022); Durmus et al. (2023); Atari et al. (2023). Since the composition of pretraining data has been shown to impact model behavior Kandpal et al. (2023); Razeghi et al. (2022); Chang et al. (2023); Gonen et al. (2023), documentation of this data facilitates informed and appropriate application of models Gebru et al. (2021).
@@
In our work, we argue that it is additionally important to examine how data is transformed prior to pretraining, and document the implications of these transformation steps. LLM pretraining data curation involves many decision points, which may be motivated by performance on popular benchmarks (e.g. Rae et al., 2021), or simply by some notion of text “quality.” There remain many under-examined assumptions within data curation pipelines, which vary subtly across models Soldaini et al. (2024); Penedo et al. (2023).
Figure 1: A paraphrased excerpt from a website’s about page, with extracted social dimensions highlighted. We use self-descriptions like this one from Common Crawl, which is frequently used as LLM pretraining data, to examine the social effects of data curation filters.
@@
We provide a new dataset and framework, AboutMe, for documenting data filtering’s effects on web text grounded in social and geographic contexts. Sociolinguistic analyses in NLP are limited by a lack of large-scale, self-reported sociodemographic information tied to language data Holstein et al. (2019); Andrus et al. (2021). Though text can be attributed to broad sources, e.g. Wikipedia, the backgrounds of content creators at more granular levels are often unknown. In particular, web crawl data lacks consistent and substantive user metadata. Our study leverages existing structure found in web data. Specifically, some websites include pages delineated to be about the website creator, such as an “about me” page (Figure 1). Thus, we are able to identify whose language is represented in web scraped text at an unprecedented scale.
@@
From websites’ about pages, we measure their topical interests, their positioning as individuals or organizations, their self-identified social roles, and their associated geographic locations (§2). We then apply ten “quality” and English ID filters drawn from prior literature on LLM development (§3) onto these websites, show whose pages are removed or retained, and investigate possible reasons for filters’ preferences (§4). Together, our experiments uncover behavioral patterns within and across filters tied to aspects of websites’ provenance. We find that model-based “quality” filters’ implicit preferences for certain topical domains lead to text specific to different roles and occupations being removed at varying rates. In addition, English content associated with non-anglophone regions of the world can be removed due to filtering approaches that assume pages are monolingual.
We release our dataset, reproduced filters, and other resources to facilitate future work:
@@ 2 Extracting Social Dimensions from about Pages
Sociolinguists conceptualize language as a performance of one’s social identity, or membership in a social group Nguyen et al. (2016). Websites’ about pages capture aspects of their creators’ social identities that they deem salient and significant enough to mention in a summary (Figure 1). Thus, these self-descriptions can help delineate meaningful differences in language varieties and use. We extract social aspects that are present across large sets of pages using automated methods, some of which we contribute as novel approaches. We do not examine attributes such as race or gender, as these are less commonly explicitly stated on about pages and may raise the risk of mismeasurement (§8).
@@
2.1 Data preprocessing
AboutMe is derived from twenty four public snapshots of Common Crawl collected between 2020–05 and 2023–06. We extract text using CCNet Wenzek et al. (2020) and deduplicate URLs across all snapshots. Our study focuses on data curation of English LLMs, and our pipeline for identifying social aspects of websites uses methods that work best for English. Thus, we limit our study to CCNet’s outputted webpages that have a fastText English ID score $>0.5$ Joulin et al. (2016b, a).
From this Common Crawl data, we identify websites that include an about page, or URL paths containing about, about-me, about-us, or bio (Appendix A). We then pair each about page with a random page on the same website. AboutMe thus contains both information about the creator/s of a website and a sample of their textual content (Table 1). Though we use this dataset to study LLM data curation practices, text linked to their creators’ self-descriptions can also facilitate research on self-presentation Sun et al. (2023); Pathak et al. (2021) and language variation Nguyen et al. (2016).
2.2 Topical interests
First, we treat about pages as summaries of website creators’ interests and topical focus. Following past work on the unsupervised discovery of domains Gururangan et al. (2023), we embed about pages using unigram counts and tf-idf Manning et al. (2008) and cluster them with balanced $k$-means. We set $k=50$ and surface a wide range of topical clusters in our data, including design, finance, food, religion, and travel (Figure 2, Appendix C). Thus, this clustering step provides an initial broad overview of who and what is in AboutMe.
@@ 2.2 Topical interests
Website creators can range from casual bloggers to larger corporations. We deem webpages with urls that contain about-me or bio as individuals, while those that are labeled as about-us are organizations. However, some pages are ambiguously labeled with about, and so we classify these into individuals or organizations by training a binary random forest classifier on labeled pages. Classifier inputs include several count features that are agnostic to pages’ topical content: the proportion of words on an about page in common pronoun series (he, she, they, we, I), the proportion of words that are tagged as a person by spaCy named-entity recognition, and the number of unique person first tokens. Our classifier achieves an average macro F1 of 89.2, via 5-fold cross validation on 10k examples per class. Hyperparameter choices, classifier confidence, and other implementation details are in Appendix D.1. By applying this classifier, we find that three-fourths of hostnames are organizations rather than individuals (Table 1).
@@ 2.4 Social roles
String-matching can be imprecise due to polysemy and mentions of other people on about pages (e.g. a customer). Thus, our role extraction approach targets explicit expressions of self-identification (e.g. I am a designer, entrepreneur, and mother). We hand-label a sample of 1K about page sentences spanning a diverse set of potential roles, [^1] https://en.wiktionary.org/w/index.php?title=Category:en:People and treat role extraction as a binary, sentence-level token classification task. Our full criteria for role annotation can be found in Appendix F.1, and we achieved high agreement (Cohen’s $\kappa$ = 0.836).
@@
With this approach, we are able to identify social roles on 77.7% of all individuals’ about pages (Table 1), and pages that have any roles contain 5.5 on average (SD = 9.9). When possible, we group terms into occupations based on a taxonomy created by the Occupational Information Network, or O*NET Peterson et al. (2001), e.g. the roles attorney and lawyer are in the occupation of Lawyer in the Legal occupation family (Table 2, Appendix F.3). For our filtering rate analysis (§4), we include 780 social roles that occur at least 1K times in AboutMe.
@@
Arts, Design, Entertainment, Sports, & Media 1.1M artist, director, designer, writer, photographer, musician, player
Production 620K designer, engineer, maker, builder, operator, mechanic
Community & Social Service 452K therapist, educator, advisor, pastor, activist, social worker
Computer & Mathematical 365K engineer, developer, scientist, strategist, programmer
Educational Instruction & Library 308K teacher, professor, lecturer, curator, tutor, graduate student
Table 2: Five most common occupation families in AboutMe, by website count, with example social roles. Additional examples can be found in Appendix F.2-F.3.
@@ 2.5 Geography
We geoparse locations on about pages using Mordecai3, which tags named locations, retrieves candidate matches from a GeoNames index, and disambiguates them using textual context Halterman (2023). For example, I’m from Alexandria, Virginia would be geoparsed to a location in the United States instead of Egypt. Mordecai3 is free and uses a local index, and so it can be scaled up to millions of webpages. With this approach, we are able to tag 63.2% of websites in AboutMe with at least one country. This coverage exceeds the 10.38% obtained by using country-specific top-level domains, e.g. .uk for the United Kingdom Cook and Brinton (2017).
Locations mentioned on a page are often associated in some way with the creator/s of a website, but the strength of this association can vary. For example, a company may have been founded in one country, but ships products to another. Through manual annotation of locations in 200 about pages, we find that 79.46% of websites with geoparsed countries originate from or reside in the most frequently referenced country on their about pages (evaluation details in Appendix G.1). Thus, we label each website with its most frequent country, yielding an overall website-level labeling accuracy of 91.0%. Due to the current global digital divide and our focus on English, the majority of websites in AboutMe are labeled with the United States and United Kingdom, with a long tail of other countries (Figure 3). For analysis, we group countries into 5 continental regions and 15 subregions delineated by the United Nations (Appendix G.2).
@@ 3 Pretraining Data Filters
Figure 3: Common continental subregions in AboutMe. The most frequent countries are the United States, United Kingdom, India, Canada, Australia, China, Germany, New Zealand, Italy, and South Africa (Appendix G).
Filter Examples of prior use Removal strategy
[…]
Raw data is transformed into pretraining data for LLMs through a variety of steps (Penedo et al., 2023; Soldaini et al., 2024, inter alia), including deduplication Lee et al. (2022), decontamination (e.g. Touvron et al., 2023b), and explicit content filtering (e.g. OpenAI, 2023). We focus on analyzing the effects of “quality” filtering and English langID. The former is motivated by the subjectivity of how quality should be defined, and the latter by ongoing uncertainty around whether langID is robust to a wide range of language varieties Seargeant and Tagg (2011); Caswell et al. (2020).
Since web text contains noisy content Eisenstein (2013), removing or downsampling “low quality” text is common in LLM development (Table 3). However, mismatch between filtering outcomes and downstream objectives can lead to performance degradation on some tasks Gao (2021); Longpre et al. (2023), or disfavor content written by minoritized populations Gururangan et al. (2022). LangID is also a common step in data curation pipelines (Table 3). It can be used at the document-level as an initial filter for language-specific (e.g. English-only) models, or to measure and adjust the composition of pretraining data for multilingual models Xue et al. (2021). However, popular langID systems are imperfect, for reasons such as training and application domain mismatch and confusion between similar languages Kreutzer et al. (2022); Caswell et al. (2020).
We critically examine ten document-level quality and English filters that are sufficiently documented in prior work (Table 3). Appendix B includes additional details on the reproduction of each filter. Descriptions of pretraining data curation are sometimes too vague or non-existent to allow for exact replication OpenAI (2023), but multiple recent and prominent LLMs still allude to the use of model- and heuristic-based data filters Touvron et al. (2023a); Gemini Team et al. (2023); Chowdhery et al. (2023).
@@
We experiment with quality filters that score text based on their similarity to some chosen “high quality” reference corpora. We name these filters based on the reference corpora used to train them: WikiWebBooks, OpenWeb, Wiki, and WikiRefs (Table 3). We use Gururangan et al. (2022)’s replication of GPT-3’s binary logistic regression quality classifier and only vary the positive “high quality” class. The negative class is a fixed set of tokens from the September 2019 dump of Common Crawl, and each class contains approximately 300M tokens. We also compare Wiki to a perplexity-based text scorer, Wikippl, which uses a 5-gram Kneser-Ney language model trained on Wikipedia instead of a classifier Wenzek et al. (2020); Laurençon et al. (2022); Muennighoff et al. (2023); Marion et al. (2023).
@@
Quality: Wiki Quality: Wikippl English: fastText
@@ 4 Whose websites are filtered?
In this section, we overview the effects of data filters on sampled webpages grouped by social dimensions identified from their about pages. Broadly, we examine the degree of consensus among filters when scoring pages, and identify themes that characterize their behavior. Within some dimensions, we also investigate whether filtering rates reflect systemic differences in power and status among social groups Blank (2013); Davis (2018).
Past LLMs have chosen a range of score cutoffs and sampling mechanisms to reduce undesirable text (Table 3). With this variation in mind, we examine the outcome of model-based filters through the lens of two contrasting scenarios. First, whose pages are least affected, or retained, if we were to keep only the documents within a top percentile of scores? Second, whose pages are most affected, or removed, if we were to filter those at a very bottom percentile? We select top and bottom cutoff percentiles of 10% and 90%, though for CLD2 and langdetect, a large number of score ties meant that that cutoffs for both scenarios were 5.2% and 8.7%, respectively. For rule-based filters, we use cutoffs specified by Rae et al. (2021). All cutoffs are listed in Appendix B.
@@
Quality: Wiki Quality: Wikippl English: fastText